OpenStudy (zab505):

In the polynomial below, what number should replace the question mark to produce a difference of squares? x2 + ?x - 25 A. 25 B. 5 C. 0 D. 10

OpenStudy (zzr0ck3r):

hint \(a^2-b^2 = (a-b)(a+b)\)

OpenStudy (zzr0ck3r):

\(x^2-25=x^2-5^2=(x-5)(x+5)\)

OpenStudy (zzr0ck3r):

so what is the middle term?

OpenStudy (zab505):

5?

OpenStudy (zzr0ck3r):

\(x^2+\color{blue}{b}x-25=x^2-25\) what does \(\color{blue}{b}\) have to be for this to be true?

OpenStudy (zzr0ck3r):

0

OpenStudy (zab505):

ahh, Thank you so much!